

********************************************************************
UARS WEEKLY STATUS REPORT - 28 OCT 2005
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004273 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005099 has been processed and is waiting for archival
at the DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004259. MEPS data between 2004260 and
2005099 has been processed and is waiting for DAAC archival.
HEPSB data files are version 1. HEPSB data files between 1991274 and
1991366 have been processed and are waiting for DAAC archival.
Processing of all PEM DAAC data files has been suspended until the
DAAC catches up with archiving all processed data files. An error was
discovered in the DAAC IP which caused problems with DAAC access of PEM
Level 2 data. This was corrected at SwRI and the DAAC was successfully
able to access SwRI. SwRI was waiting for the DAAC to catch up on Level 2
PEM files. However, during this process, the DAAC again ran into access
problems. It was found that the DAAC IP node had changed again which
disrupted DAAC access of PEM Level 2 data. Currently the DAAC and PEM are
investigating the reason for another change in the DAAC IP. A solution is
pending.
Checks on the PEM data availability through the DAAC have been conducted
by SwRI. These tests were prompted by a user of PEM data who has had trouble
accessing the PEM data in the DAAC. The documentation link to the PEM HEPSA
Level 2 Document still points to the old version file pem_l2_hepsa_v01.txt
instead of the v02 file for hepsa data. Also, there is no MEPS entry for the
MEPS data set. These problems have been reported to help@daac.gsfc.nasa.gov
in mid-October.
Tests of the AXIS1 data file formats have been successful. The AXIS1
format design has been coded. A test file has been generated. A check
program with processing scripts have also been generated. Example binary
read program is complete and verified. These results have been sent to
the project. Creation of the AXIS1 data documents for DAAC data will
begin shortly.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG is participating in the on-going UARS end of mission
planning activities.
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 15 October 2005 (Orbit 77153, GMT Day 288) through
21 October 2005 (Orbit 77260, GMT Day 294).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
-NONE THIS REPORT PERIOD
The beta angle ranged between the angles of 38.4 degrees to 15.1
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of -2 degrees on 24 October 2005 (DOY 297).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. TMON 45 and 46 are
maintaining the relay operations for battery 2. Battery 2 and Battery
3 are currently maintaining greater than 23.6 V end-of-night (EON)
load bus voltage. Battery temperatures are stable with a temperature
delta between Battery 2 and Battery 3 of 2.56 to 2.23 degrees C.
Additional battery performance data for each battery on 14 October
2005 and 21 October 2005 are provided in Appendix D.
The clock error ranged from 11.7 to -10.6 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 07
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continues to closely monitor the condition of
battery 2 and battery 3. Action was taken on the continual uneven
load sharing between battery 2 and 3. It was determined that the PMON
calculated C/D ratio for battery 2 was not accurate. The reason for
the inaccuracy is still not completely understood. Manual calculation
of the battery C/D ratio using a full power orbital cycle, provided by
way of tape recorder dumps recorded during the orbit lowering burns,
gave the FOT probable cause that the C/D ratio calculation by PMON was
anywhere from 3% to 6% off. The FOT believes that this discrepancy
started about on 12 September 2005 when the FOT changed TMON 45 and
46. The consequence of undercharging battery 2 was that battery 3
carried the bulk of the spacecraft power load.
To remedy this undercharging of battery 2 the FOT and power branch
made minor tweaks to TMON 45 and it's timing on 17 October 2005. The
amount of time that the TMON waited for to take battery 2 off relay
was increased from 32 seconds to 96 seconds. This allowed roughly one
minute of "overcharge" to go into battery 2. The results of this
change were recognized almost immediately. Within 3 orbits of the
change, the DOD and temperature on battery 2 began increasing. Within
3 days of the change, the FOT reduced the VT charging level from 6 to
5 since the batteries were performing much better. Currently the
battery load sharing has been much better. Battery 2 supplies almost
1/3 of the power load while battery 3 supplies the remaining amount.
This improved load sharing bodes well for the upcoming instrument turn
on after UARS yaws to the forward flight direction. NASA power branch
has been providing daily support in monitoring the UARS battery
condition. The FOT has been authorized to disable all platform and
instrument heaters on the SSPP. This will be done at the earliest
available opportunity.
The FOT successfully completed both planned orbit-lowering burns this
week. The orbit lowering burns on 18 October 2005 and 20 October 2005
went as expected. Both burns started as planned at 291/1443 and
293/1209. Each ran for 18 minutes. The calculated duty cycle for the
burn on the 18th was 77.4%, and the duty cycle for the burn on the
20th was 77.3%. FDF calculated that perigee was lowered by 27.1 km at
the completion of the burn on the 18th, and then lowered by an
additional 25.5 km at the completion of the burn on the 20th. The now
elliptical orbit of UARS is approximately 557 x 383 KM.
MEETINGS
The FOT attended a CAM meeting on 17 October 2005. The orbit lowering
burn script and procedures for the burn on 18 October 2005 were
reviewed and approved during this meeting. FDF also presented their
predicted pre burn position and velocity for the upcoming burn.
The FOT attended a combined debrief and CAM meeting on 19 October
2005. Results of the burn on 18 October 2005 were presented by FDF
and the FOT. The orbit lowering burn script and procedures for the
burn on 20 October 2005 were reviewed and approved during this
meeting. FDF also presented their predicted pre burn position and
velocity for the upcoming burn.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 24 October 2005 (DOY 297).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 21 October 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
77196 290/1927 05-370 OBC TMON 45 - CHANGED OOL COUNTS
77199 290/2336 05-371 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
77208 291/1305 05-372 MACS ORBIT ADJUST BURN #4 - STEP 1
77208 291/1334 05-372 MACS ORBIT ADJUST BURN #4 - STEP 2-11
77208 291/1421 05-372 MACS ORBIT ADJUST BURN #4 - STEP 12
77208 291/1439 05-372 MACS ORBIT ADJUST BURN #4 - STEP 13-23
77209 291/1519 05-372 MACS ORBIT ADJUST BURN #4 - STEP 25-27
77209 291/1627 05-372 MACS ORBIT ADJUST BURN #4 - STEP 24
77217 292/0422 05-373 OBC CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
77224 292/1615 05-374 OBC EPHEM WINOPEN / WINDEF
77236 293/1019 05-375 MACS ORBIT ADJUST BURN #5 - STEP 1
77236 293/1059 05-375 MACS ORBIT ADJUST BURN #5 - STEP 2-11
77237 293/1201 05-375 MACS ORBIT ADJUST BURN #5 - STEP 12-22
77237 293/1241 05-375 MACS ORBIT ADJUST BURN #5 - STEP 23,25-27
77238 293/1346 05-376 MPS/TMON TMON 46 RELAY TIME / VT LEVEL 5
77239 293/1426 05-375 MACS ORBIT ADJUST BURN #5 - STEP 24
77240 293/1631 05-377 MACS EPHEM WINDEF
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 21 October 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 21 October 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
77199 290/2339 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
77217 292/0430 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 21 October 2005
14 October 2005 - GMT Day 287 Beta = 41.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -1.48/ -2.42 +00.0/+00.0 04.6 0.00 0.0
2 -700.0/-700.0 -0.54/ -1.79 +13.6/-06.8 25.4 1.05 3.3
3 011.2/-134.4 +10.40/ +8.40 +44.4/-25.6 25.4 1.04 25.8
21 October 2005 - GMT Day 294 Beta = 15.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -2.42/ -3.68 +00.0/+00.0 04.6 0.00 0.0
2 -700.0/-700.0 +3.23/ +2.60 +23.2/-12.0 26.1 1.06 11.1
3 224.0/ 00.0 +5.79/ +4.83 +31.2/-22.8 25.8 1.02 22.9
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
77238 293/1346 LV6 -> LV5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 21 October 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-047 FOT 77157 288/0551 30832 Generic Late Acquisition
05-047 FOT 77211 291/1944 30832 Generic Late Acquisition
05-052 FOT 77237 293/1241 41080 Post-burn late Acq-2:43
recov
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-049 FOT 77094 284/0157 41076 WSGT unable to support
05-050 FOT 77100 284/1213 41080 Neg Acquisition
05-051 FOT 77101 284/1345 41080 Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-042 MPS 76339 233/2347 N/A BATTERY 2 DIFFV -700mV
05-047 FOT See List Oct 05 30832 Generic Late Acquisition
05-052 FOT 77237 293/1241 41080 Post-burn late Acq-2:43
recov